Italian Queue 2014 Prev Next

Jun 21-22

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: O. Khun
Instrument: MODS/LBC

Summary

At enclosure opening the sky is clear with no wind, and 34% humidity
At the beginning of sxcientific operations the sky is clear, whit 1.0 seeing
Stuck of the readout in the MODS red channel. The SA sees the image in the mods computer, but it is not written in /newdata. Aferter one out of efforts spent in calling peaople and leaving them messages, we ask to switch instrument.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Elias_Rosa 11FE dual grating 1h Data acquired, but still don't saved at the end of the night. LOST the day after
Gilli4_A12 J1641 r-Sloan, i-sloan 1.5h, 1.5h partly within seen spec
Standard Sa112 r-Sloan, i-sloan, z-Sloan   airmas 1.18
Gilli4_A14 J2229 r-Sloan, i-sloan 30m, 30m within seeing specs, the last 4 frames with increasing backgorund

Details

02:20 Opening the enclosure

02.41 We start running calibrations for LBC
  • Run SkyFlat _rr_r1502d+29_pa0_5.ob
  • Run SkyFlat _rr_r1502d+29_pa180_5.ob
  • Run SkyFlat _gi_r1502d+29_pa180_5.ob
  • Run SkyFlat _gi_r1502d+29_pa0_5.ob

03.14 We ask the TO to switch to MODS

03:35 We are authorized to use MODS and slew to Elias_Rosa_47_DDT

03:37 acqMODS SN2011fe_ls.acq mods1r.0001-0003
  • Stuck of the readout in the red channel. The SA restart the red channel computer.
  • modsAlign -l mods1r.20140622.0001.fits mods1r.20140622.0002.fits

04:00 execMODS SN2011fe_ls.obs mods1b.0002-000 mods1r.0004-000
  • once at the end, the read out is stuck. The SA sees the image on the MODS computer but it not saved in /newdata. This is 1h integration image. We try to call, but no one is answering. No one answer after 1h of efforts (at least the seeing is bad.... >1.5). Since the seeing is bad, we ask the TO to switch to LUCI.

06:10 Problem with the guide camera

06:36 Temperature warning from the water used to refrigerate the telescope. Mike is going to give a look

06:28 we are authorized to use LUCI
  • executeLUCIscript.sh 7C1756Mask1_MOS_210_HeIJ_tell.acq luci.20140622.0001-0002
  • the probe got stuck. Since the seeing now is good (< 0.8) we ask the TO to switch to LBC

** We are authorized to use LBC

  • Gilli

07:02 Which to Gilli 12
  • dofpia,/X2,/first
  • Lot of coma. Increase Z8 by hand (+2000)
  • dofpia,/X2

07:16 pointing_A12.ob
  • Error in xpa communication between ds9 and lbc_rangebal
  • Run ds9 from IRAF
  • run lbc_rangebal again

07:24 ri_A12_1.ob
  • error from the filter weehl. Script crashed. The first blue image is missed. We run again the script from the beginning
  • variable seeing: 0.9-1.2

08.10 Run dofpia. /X2 on the scientific field

08:14 Run ri_A12_2.ob
  • Error in the blue channel. After 5th exposure (3 red, 2 blue)
  • Run the ob again
  • mostly out of seeing specs

09.05 dofpia. /X2 on the scientific field Gilli4_A12
  • Run ri_A12_3.ob

09:57 Run SA112_r_iz.ob

10:10 Slew to Gilli4_A14 field
  • run dofpia on the pointing ob
  • lbcrangebal B=101633 R=101623

10:20 ri_A14_1.ob
  • within specs, seeing 0.8 in R and 1.0 in B
  • last 4 offsets with increasing background --> to be repeated

11:00 We start running sky flats
  • run SkyFlat _gz_r2258d00_pa0_5.ob a bit low in z-band
  • run SkyFlat _gz_r2258d00_pa180_5.ob saturated in g, good in z
  • run SkyFlat _gz_r2258d00_pa0_5.ob scale=0.5

11:49 Enclosure closing
Topic revision: r7 - 23 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback